Solve the following system of equations by any method:
4x + 3y = 12
y = x – 10

Answer by TimothyLamb(4379)

4x + 3y = 12
y = x - 10
---
put the system of linear equations into standard form
---
4x + 3y = 12
x - y = 10
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 6
y = -4
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---